OpenStudy (anonymous):

find the slope of each line 2x+3y=14

OpenStudy (anonymous):

Slope of a line ax+by=c is -a/b

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

WHAT U MEAN

OpenStudy (anonymous):

Compare ax+by=c with 2x+3y=14 and tell me the values of a and b

OpenStudy (anonymous):

DO I ADD THEM UP

OpenStudy (anonymous):

Nope u need to find -a/b

OpenStudy (anonymous):

DO I DIVIDE OR CAN U GUIDE ME THROUGH THE STEPS

OpenStudy (anonymous):

Just compare the two equations and u will get: a=2 AND b=3 Thus, The slope of the line =-a/b = -2/3

OpenStudy (anonymous):

SO WHAT SHOULD I PUT AS MY ANSWER

OpenStudy (anonymous):

-2/3